# A Unifying Framework for Parallelizing Sequential Models with Linear Dynamical Systems

### v1

@@ B.1 A very simple example: multiplying matrices
The parallel scan (Stone, 1973; Blelloch, 1990), also known as the associative scan and, colloquially, pscan, is a well-known primitive in the parallel computing literature (Hillis & Steele Jr, 1986; Ladner & Fischer, 1980; Lakshmivarahan & Dhall, 1994). The core idea of the parallel scan is a divide-and-conquer algorithm. We illustrate this point in the simple example of multiplying a series of matrices together.
@@
A core insight of the parallel scan is that matrix multiplication is closed; that is, if $A_{s}\in\mathbb{R}^{D\times D}$ and $A_{t}\in\mathbb{R}^{D\times D}$, then $A_{t}A_{s}\in\mathbb{R}^{D\times D}$. Thus, matrix products can be computed recursively in pairs, as illustrated in Figure˜4.
@@ B.2 Detail #1: Parallel scans for arbitrary binary associative operators
Matrix multiplication is an associative operator, as $A_{3}(A_{2}A_{1})=(A_{3}A_{2})A_{1}$. In general, consider a binary associative operator $\otimes$, which would satisfy $q_{3}\otimes(q_{2}\otimes q_{1})=(q_{3}\otimes q_{2})\otimes q_{1}$. Now, let us further assume that this binary associative operator closed:
@@
    This observation that composition of affine functions can be parallelized with the associative scan is what lets us parallelize LDSs. The parallelization of LDSs is what allows for parallel computation in many important architectures in sequence modeling, such as linear RNNs (Martin & Cundy, 2018; Orvieto et al., 2023), deep SSMs (Smith et al., 2023; Gu & Dao, 2024), and nonlinear [^5] The parallel scan is used in nonlinear RNNs via the iterative fixed-point methods discussed in this paper, i.e. Newton and quasi-Newton iterations, see Algorithm 1. RNNs (Lim et al., 2024; Gonzalez et al., 2024; Farsang et al., 2025). This parallel scan for LDSs is the core primitive uniting the fixed-point methods discussed in this paper.
  • Kalman filtering and smoothing: Parallel scans can also be utilized in probabilistic modeling. A standard probabilistic model is the linear Gaussian state space model (LGSSM), where the latent variables $x_{t}$ follow linear dynamics with Gaussian noise, and emit observations $y_{t}$ with linear readouts with Gaussian noise (Murphy, 2023; Särkkä & Svensson, 2023). See Figure˜5.
@@
    Two canonical inferential targets in the LGSSM are the filtering distributions, $p(x_{t}\mid y_{1:t})$, and the smoothing distributions, $p(x_{t}\mid y_{1:T})$. The Kalman filter (Kalman, 1960) and Rauch-Tung-Striebel (RTS) smoother (Rauch et al., 1965) obtain the filtering and smoothing distributions (respectively) in an LGSSM. The Kalman filter makes a single pass forward in time to get the filtering distributions, while the RTS smoother then makes an additional pass backwards in time to get the smoothing distributions. Both the Kalman filter and RTS smoother would seem to be inherently sequential algorithms, requiring $\mathcal{O}(T)$ time. However, Särkkä & García-Fernández (2021) demonstrated that the Kalman filter and RTS smoother can also be parallelized over the sequence length via the construction of custom binary associative operators and a parallel scan. While we leave the details of this construction to Särkkä & García-Fernández (2021), we note that it is intuitively plausible to be able to parallelize filtering and smoothing in an LGSSM with a parallel scan because
@@
    These parallel filtering and smoothing algorithms are useful in machine learning, allowing for parallelization of structured variational autoencoders (Johnson et al., 2016; Zhao & Linderman, 2023). Similar approaches also work for Hidden Markov Models (Hassan et al., 2021) and for computing log-normalizing constants (Hu et al., 2025).
  • Parallelizing fixed point iterations: Finally, these parallel filtering and smoothing are directly applicable to the types of parallel fixed-point iterations that are the focus of this paper. In particular, the ELK algorithm (Gonzalez et al., 2024)Evaluating Levenberg-Marquardt via Kalman—stabilizes the Newton iterations (Lim et al., 2024) we discuss in this paper using the Levenberg-Marquardt (trust-region) method (Levenberg, 1944; Marquardt, 1963). The trust-region updates are able to be computed using a parallel scan because they are equivalent to a Kalman smoother in an appropriately constructed LGSSM (Särkkä & Svensson, 2020).
@@ B.4 Implementation considerations
This gentle introduction provides the main ideas and intuition of the parallel scan: closed binary associative operators leading a divide-and-conquer algorithm to leverage parallel processors to compute sequential compositions in sublinear time. However, there are also a host of implementation details for using the parallel scan when programming on accelerated hardware like GPUs (Harris et al., 2007; Yang et al., 2024; Sarnthein, 2025). For example, the presence of a general-purpose parallel scan is, as of the time of writing, a major difference between JAX (Bradbury et al., 2018) and PyTorch (Paszke et al., 2019), two leading Python libraries for deep learning. JAX has a general purpose parallel scan (jax.lax.associative_scan) as a fundamental primitive, which allows for implementation of a wide range of parallel scans. For example, dynamax, a JAX library for probabilistic state space modeling (Linderman et al., 2025), implements the parallel filtering and smoothing algorithms from Särkkä & García-Fernández (2021). In contrast, PyTorch currently has only torch.cumsum, which is the parallel scan where the binary associative operator is addition [^8] Although Heinsen (2023) shows that clever uses of torch.cumsum can parallelize scalar/diagonal LDSs, of the type that are used in quasi-Newton iterations (eq. 6)., and torch.cumprod (for scalar multiplication). This difference is why we implement the experiments in this paper in JAX. This lack of a general purpose parallel scan in PyTorch has also led to the custom development of highly-optimized, hardware-aware custom CUDA kernels for parallel scans, such as in Mamba (Gu & Dao, 2024), a leading SSM for language modeling. There also exist useful implementations of parallel scans for scalar/diagonal LDSs in PyTorch such as Kyrylov (2024), which we used to implement quasi-Newton iterations in PyTorch in this repo: https://github.com/lindermanlab/elk-torch
Appendix C Experimental Details and Additional Experiments

### v2

@@ B.1 A very simple example: multiplying matrices
The parallel scan (Stone, 1973; Blelloch, 1990), also known as the associative scan and, colloquially, pscan, is a well-known primitive in the parallel computing literature (Hillis and Steele Jr, 1986; Ladner and Fischer, 1980; Lakshmivarahan and Dhall, 1994). The core idea of the parallel scan is a divide-and-conquer algorithm. We illustrate this point in the simple example of multiplying a series of matrices together.
@@
A core insight of the parallel scan is that matrix multiplication is closed; that is, if $A_{s}\in\mathbb{R}^{D\times D}$ and $A_{t}\in\mathbb{R}^{D\times D}$, then $A_{t}A_{s}\in\mathbb{R}^{D\times D}$. Thus, matrix products can be computed recursively in pairs, as illustrated in Figure˜5.
@@ B.2 Detail #1: Parallel scans for arbitrary binary associative operators
Matrix multiplication is an associative operator, as $A_{3}(A_{2}A_{1})=(A_{3}A_{2})A_{1}$. In general, consider a binary associative operator $\otimes$, which would satisfy $q_{3}\otimes(q_{2}\otimes q_{1})=(q_{3}\otimes q_{2})\otimes q_{1}$. Now, let us further assume that this binary associative operator is closed:
@@
    This observation that composition of affine functions can be parallelized with the associative scan is what lets us parallelize LDSs. The parallelization of LDSs is what allows for parallel computation in many important architectures in sequence modeling, such as linear RNNs (Martin and Cundy, 2018; Orvieto et al., 2023), deep SSMs (Smith et al., 2023; Gu and Dao, 2024), and nonlinear [^6] The parallel scan is used in nonlinear RNNs via the iterative fixed-point methods discussed in this paper, i.e. Newton and quasi-Newton iterations, see Algorithm 1. RNNs (Lim et al., 2024; Gonzalez et al., 2024; Farsang and Grosu, 2025; Danieli et al., 2026). This parallel scan for LDSs is the core primitive uniting the fixed-point methods discussed in this paper.
  • Kalman filtering and smoothing: Parallel scans can also be utilized in probabilistic modeling. A standard probabilistic model is the linear Gaussian state space model (LGSSM), where the latent variables $x_{t}$ follow linear dynamics with Gaussian noise, and emit observations $y_{t}$ with linear readouts with Gaussian noise (Murphy, 2023; Särkkä and Svensson, 2023). See Figure˜6.
@@
    Two canonical inferential targets in the LGSSM are the filtering distributions, $p(x_{t}\mid y_{1:t})$, and the smoothing distributions, $p(x_{t}\mid y_{1:T})$. The Kalman filter (Kalman, 1960) and Rauch-Tung-Striebel (RTS) smoother (Rauch et al., 1965) obtain the filtering and smoothing distributions (respectively) in an LGSSM. The Kalman filter makes a single pass forward in time to get the filtering distributions, while the RTS smoother then makes an additional pass backwards in time to get the smoothing distributions. Both the Kalman filter and RTS smoother would seem to be inherently sequential algorithms, requiring $\mathcal{O}(T)$ time. However, Särkkä and García-Fernández (2021) demonstrated that the Kalman filter and RTS smoother can also be parallelized over the sequence length via the construction of custom binary associative operators and a parallel scan. While we leave the details of this construction to Särkkä and García-Fernández (2021), we note that it is intuitively plausible to be able to parallelize filtering and smoothing in an LGSSM with a parallel scan because
@@
    These parallel filtering and smoothing algorithms are useful in machine learning, allowing for parallelization of structured variational autoencoders (Johnson et al., 2016; Zhao and Linderman, 2023). Similar approaches also work for Hidden Markov Models (Hassan et al., 2021) and for computing log-normalizing constants (Hu et al., 2025).
  • Parallelizing fixed point iterations: Finally, these parallel filtering and smoothing are directly applicable to the types of parallel fixed-point iterations that are the focus of this paper. In particular, the ELK algorithm (Gonzalez et al., 2024)Evaluating Levenberg-Marquardt via Kalman—stabilizes the Newton iterations (Lim et al., 2024) we discuss in this paper using the Levenberg-Marquardt (trust-region) method (Levenberg, 1944; Marquardt, 1963). The trust-region updates are able to be computed using a parallel scan because they are equivalent to a Kalman smoother in an appropriately constructed LGSSM (Särkkä and Svensson, 2020).
@@ B.4 Implementation considerations
This gentle introduction provides the main ideas and intuition of the parallel scan: closed binary associative operators leading a divide-and-conquer algorithm to leverage parallel processors to compute sequential compositions in sublinear time. However, there are also a host of implementation details for using the parallel scan when programming on accelerated hardware like GPUs (Harris et al., 2007; Yang et al., 2024; Sarnthein, 2025). For example, the presence of a general-purpose parallel scan is, as of the time of writing, a major difference between JAX (Bradbury et al., 2018) and PyTorch (Paszke et al., 2019), two leading Python libraries for deep learning. JAX has a general purpose parallel scan (jax.lax.associative_scan) as a fundamental primitive, which allows for implementation of a wide range of parallel scans. For example, dynamax, a JAX library for probabilistic state space modeling (Linderman et al., 2025), implements the parallel filtering and smoothing algorithms from Särkkä and García-Fernández (2021). In contrast, PyTorch currently has only torch.cumsum, which is the parallel scan where the binary associative operator is addition [^9] Although Heinsen (2023) shows that clever uses of torch.cumsum can parallelize scalar/diagonal LDSs, of the type that are used in quasi-Newton iterations (eq. 6)., and torch.cumprod (for scalar multiplication). This difference is why we implement the experiments in this paper in JAX. This lack of a general purpose parallel scan in PyTorch has also led to the custom development of highly-optimized, hardware-aware custom CUDA kernels for parallel scans. These custom parallel scans appear most prominently in Mamba (Gu and Dao, 2024), a leading SSM for language modeling, and ParaRNN (Danieli et al., 2026), which applies the Newton iterations discussed in this paper to 7B parameter nonlinear RNNs to achieve strong language modeling performance. There also exist useful implementations of parallel scans for scalar/diagonal LDSs in PyTorch such as Kyrylov (2024), which we used to implement quasi-Newton iterations in PyTorch in this repo: https://github.com/lindermanlab/elk-torch
Appendix C Further Discussion of Convergence Analysis
In this appendix, we provide a deeper discussion of the details, intuitions, and limitations of the convergence analysis presented in Section˜3.
C.1 Formula for approximate Jacobian
For completeness, we provide the formula for the block-bidiagonal approximate Jacobian $\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})\in\mathbb{R}^{TD\times TD}$ that the fixed-point methods considered in this paper give rise to.
$\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}):=\begin{pmatrix}I_{D}&0&0&\ldots&0&0\\ -\tilde{A}_{2}(x_{1})&I_{D}&0&\ldots&0&0\\ 0&-\tilde{A}_{3}(x_{2})&I_{D}&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&I_{D}&0\\ 0&0&0&\ldots&-\tilde{A}_{T}(x_{T-1})&I_{D}\\ \end{pmatrix}$. (19)
For Newton’s method $\mathcal{A}_{N}$, the resulting $\widetilde{\mathbf{J}}_{N}(\mathbf{x}_{1:T})=\mathbf{J}(\mathbf{x}_{1:T})$, where $\mathbf{J}(\mathbf{x}_{1:T})$ is defined in eq.˜11. For Picard iterations, $\widetilde{\mathbf{J}}_{P}(\mathbf{x}_{1:T})$ takes the form
$\widetilde{\mathbf{J}}_{P}(\mathbf{x}_{1:T})=\begin{pmatrix}I_{D}&0&0&\ldots&0&0\\ -I_{D}&I_{D}&0&\ldots&0&0\\ 0&-I_{D}&I_{D}&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&I_{D}&0\\ 0&0&0&\ldots&-I_{D}&I_{D}\\ \end{pmatrix}$. (20)
For Jacobi iterations, $\widetilde{\mathbf{J}}_{J}(\mathbf{x}_{1:T})$ is always the identity matrix $I_{TD}$.
C.2 Limitations of Proposition˜3
Proposition˜3 only guarantees a decrease in the error when the iterate $\mathbf{x}_{1:T}^{(i)}$ is already in a basin of decrease $\mathcal{B}_{D}$ given by
$\mathcal{B}_{D}:=\left\{\mathbf{x}_{1:T}:\|\mathbf{e}(\mathbf{x}_{1:T})\|_{2}\leq 2\cdot\frac{1-\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})^{-1}\right\|_{2}\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})-\mathbf{J}(\mathbf{x}_{1:T})\right\|_{2}}{L\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})^{-1}\right\|_{2}}\right\}.$
However, since we know from Proposition 1 of Gonzalez et al. (2024) that all the fixed-point algorithms considered in this paper must eventually converge, we know that the iterates $\mathbf{x}_{1:T}^{(i)}$ must all eventually enter this basin of decrease $\mathcal{B}_{D}$ if $\mathcal{B}_{D}\neq\emptyset$. For this reason, Proposition˜3 provides helpful intuition about which fixed-point algorithms are useful for which dynamical systems.
For example, let us define the basin of linear rate $\mathcal{B}_{L}$ to comprise those $\mathbf{x}_{1:T}$ where $\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})-\mathbf{J}(\mathbf{x}_{1:T})\right\|_{2}\|\mathbf{e}^{(i)}_{1:T}\|_{2}>\frac{L}{2}(\|\mathbf{e}^{(i)}_{1:T}\|_{2}^{2})$, i.e. the expression linear in $\|\mathbf{e}^{(i)}_{1:T}\|_{2}$ on right side of equation 13 dominates the expression quadratic in $\|\mathbf{e}^{(i)}_{1:T}\|_{2}$. It follows that $\mathcal{B}_{L}$ is given by
$\mathcal{B}_{L}:=\left\{\mathbf{x}_{1:T}:\|\mathbf{e}(\mathbf{x}_{1:T})\|_{2}\leq\frac{2\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})-\mathbf{J}(\mathbf{x}_{1:T})\right\|_{2}}{L}\right\}.$
Therefore, when $\mathbf{x}_{1:T}^{(i)}\in\mathcal{B}_{D}\cap\mathcal{B}_{L}$, it follows that the norm of the error is guaranteed to decrease by a factor of $2\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})^{-1}\right\|_{2}\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})-\mathbf{J}(\mathbf{x}_{1:T}^{(i)})\right\|_{2}$. Moreover, as $\|\mathbf{e}^{(i)}_{1:T}\|_{2}$ approaches zero, the guaranteed factor of decrease approaches the value given by eq.˜15.
C.3 Extended discussion about intuitions from the rate of convergence
In this appendix, we elaborate on the intuitions about the rate of convergence presented in Section˜3.3.
C.3.1 Intuitions from Jacobian approximation error
First, elaborating on the section discussing $\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})-\mathbf{J}(\mathbf{x}_{1:T})\right\|_{2}$, we provide a proof of Lemma˜1.
Proof of Lemma˜1.
[…]
$\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})-\mathbf{J}(\mathbf{x}_{1:T})=\begin{pmatrix}0&0&0&\ldots&0&0\\ E_{2}&0&0&\ldots&0&0\\ 0&E_{3}&0&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&0&0\\ 0&0&0&\ldots&E_{T}&0\\ \end{pmatrix}$.
The spectral norm of a matrix $M$ is equal to the square root of the largest eigenvalue of $M^{\top}M$. Defining $M:=\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})-\mathbf{J}(\mathbf{x}_{1:T})$, we see that
$M^{\top}M=\begin{pmatrix}0&0&0&\ldots&0&0\\ 0&E_{2}^{\top}E_{2}&0&\ldots&0&0\\ 0&0&E_{3}^{\top}E_{3}&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&E_{T-1}^{\top}E_{T-1}&0\\ 0&0&0&\ldots&0&E_{T}^{\top}E_{T}\\ \end{pmatrix}$.
Since $M^{\top}M$ is a block-diagonal matrix, its eigenvalues are equal to the union of the eigenvalues of each of the blocks $E_{t}^{\top}E_{t}$. Thus, it follows that the maximum eigenvalue of $M^{\top}M$ is equal to the maximum of all the eigenvalues of all the matrices $E_{t}^{\top}E_{t}$, and so the maximum singular value of $\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})-\mathbf{J}(\mathbf{x}_{1:T})$ is given by $\max_{1\leq t\leq T-1}\left\|\tilde{A}_{t+1}(x_{t})-A_{t+1}(x_{t})\right\|_{2}$. ∎
C.3.2 Intuitions from the norm of the inverse of the approximate Jacobian
Next, we elaborate on our bounds for $\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})^{-1}\right\|_{2}$.
Recall that $\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})^{-1}$ takes the form (Dao and Gu, 2024; Gonzalez et al., 2025)
$\widetilde{\mathbf{J}}(\mathbf{x}_{1:4})^{-1}=\begin{pmatrix}I_{D}&0&0&0\\ \tilde{A}_{2}&I_{D}&0&0\\ \tilde{A}_{3}\tilde{A}_{2}&\tilde{A}_{3}&I_{D}&0\\ \tilde{A}_{4}\tilde{A}_{3}\tilde{A}_{2}&\tilde{A}_{4}\tilde{A}_{3}&\tilde{A}_{4}&I_{D}\end{pmatrix}$,
shown above for $T=4$.
Theorem 2 of Gonzalez et al. (2025) controls $\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})^{-1}\right\|_{2}$ in terms of $\overline{\rho}:=\sup_{2\leq t\leq T,x\in\mathbb{R}^{D}}\|\tilde{A}_{t}(x)\|_{2}$ and $\underline{\rho}:=\inf_{2\leq t\leq T,x\in\mathbb{R}^{D}}\|\tilde{A}_{t}(x)\|_{2}$ as
$\max(1,\underline{\rho}^{T-1})\leq\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})^{-1}\|_{2}\leq\frac{1-\overline{\rho}^{T}}{1-\overline{\rho}},$
for $\overline{\rho}\neq 1$. We would therefore expect that fixed-point methods that give rise to unstable LDSs with transition matrices having spectral norms much larger than one should have slower rates of convergence. Methods that give rise to unstable LDSs suffer from numerical blowup, especially for large $T$.
Moreover, in the special cases of Jacobi and Picard iterations, we can compute $\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T}^{(i)})^{-1}\right\|_{2}$ analytically. For Jacobi iterations, $\left\|\widetilde{\mathbf{J}}_{J}\right\|_{2}=1$. For Picard iterations, the expression for $\left\|\widetilde{\mathbf{J}}_{P}\right\|_{2}$ is more complicated, but it scales as $O(T)$ (see Lemma˜2 in Section˜D.1).
Because $\|\widetilde{\mathbf{J}}_{P}^{-1}\|_{2}>\|\widetilde{\mathbf{J}}_{J}^{-1}\|$ for large $T$, the formula for $\gamma$ given by eq.˜15 yields the following expectation:
In settings where the $A_{t+1}$ from Picard vs. Jacobi iterations approximates the true dynamics Jacobian $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$ equally well, we expect Jacobi iterations to converge more quickly because $\|\widetilde{\mathbf{J}}_{J}^{-1}\|<\|\widetilde{\mathbf{J}}_{P}^{-1}\|_{2}$.
We test this hypothesis in the next section with a simple simulation designed to show how Proposition˜3 provides helpful intuition about the convergence rates of different fixed-point methods.
C.4 A simulation using Proposition˜3 to distinguish between Jacobi and Picard iterations
Figure 7: Comparing Picard and Jacobi iterations on a diagonal LDS. For the underlying dynamical system $x_{t+1}=\alpha x_{t}$ with $T=100$, we plot the norm of the error $\mathbf{e}^{(i)}_{1:T}$ for Jacobi (Left) and Picard iterations (Center). In the (Right) panel, we also show the number of iterations needed for the norm of the error to go below $1\text{\times}{10}^{-5}$.
We demonstrate the helpfulness of the intuitions stemming from Proposition˜3 in a simple simulation. We consider the LDS $x_{t+1}=\alpha x_{t}$, for $x_{t}\in\mathbb{R}^{2}$. Because this is an LDS with diagonal dynamics, both the Newton and quasi-Newton iterations considered in this paper converge in one iteration. However, this simulation is useful for comparing Jacobi versus Picard iterations. This comparison is particularly fruitful in light of the formula for $\gamma$ given by eq.˜15 and Lemma˜1 because, in this setting,
$\|\widetilde{\mathbf{J}}_{J}-\mathbf{J}\|_{2}$ $=\alpha$
$\|\widetilde{\mathbf{J}}_{P}-\mathbf{J}\|_{2}$ $=1-\alpha.$
However, $\|\widetilde{\mathbf{J}}_{J}^{-1}\|_{2}=1$, while $\|\widetilde{\mathbf{J}}_{P}^{-1}\|_{2}$ scales linearly with $T$. Therefore, when comparing the number of Jacobi iterations needed to converge when the dynamics are multiplication by $\alpha$ to the number of Picard iterations needed to converge when the dynamics are multiplication by $1-\alpha$, we expect fewer Jacobi iterations should be needed than Picard iterations, as $\gamma_{J}<\gamma_{P}$.
We observe precisely this behavior in Figure˜7. For $\alpha=0.5$, when $\|\widetilde{\mathbf{J}}_{J}-\mathbf{J}\|_{2}=\|\widetilde{\mathbf{J}}_{P}-\mathbf{J}\|_{2}$ , we see that Jacobi iterations converge in far fewer iterations than Picard iterations. Moreover, when comparing the behavior of Jacobi for simulating $f_{t+1}(x_{t})=\alpha x_{t}$ with Picard for simulating $f_{t+1}(x_{t})=(1-\alpha)x_{t}$, we observe that Jacobi iterations always converge faster. However, when comparing for the same value of $\alpha$, we see that Picard can be faster than Jacobi when $\alpha$ is closer to one. This behavior makes sense, because in those settings the true Jacobian $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$ is closer to $I_{D}$ than to $\mathbf{0}$.
Moreover, we observe that in this setting, the error $\mathbf{e}^{(i)}_{1:T}$ for Jacobi iteration shows a clear linear convergence rate, as predicted by Proposition˜3. The slope of the norm of the errors of the Jacobi iterates should be $\log_{10}(\alpha)$ by eq.˜15 and Lemma˜1, and in fact those values are exactly the slopes of the lines in Figure˜7 (Left panel).
Appendix D Theoretical Details: Additional Proofs
D.1 Lemma˜2
Lemma 2.
Let $\widetilde{\mathbf{J}}_{P}$ be as in eq.˜20. Then
$\|\widetilde{\mathbf{J}}_{P}^{-1}\|_{2}=\frac{1}{2\sin\left(\frac{\pi}{2(2T+1)}\right)}$
By the small-angle approximation for sine, $\|\widetilde{\mathbf{J}}_{P}^{-1}\|_{2}$ scales as $\mathcal{O}(T)$.
Proof.
Consider
$K:=\widetilde{\mathbf{J}}_{P}^{-\top}\widetilde{\mathbf{J}}_{P}^{-1}=\begin{pmatrix}I_{D}&I_{D}&I_{D}&\ldots&I_{D}\\ I_{D}&2I_{D}&2I_{D}&\ldots&2I_{D}\\ I_{D}&2I_{D}&3I_{D}&\ldots&3I_{D}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ I_{D}&2I_{D}&3I_{D}&\ldots&TI_{D}\end{pmatrix}$.
We know that $\lambda_{\max}(K)^{1/2}=\|\widetilde{\mathbf{J}}_{P}^{-1}\|_{2}$. Since $K$ is a Kronecker product $M\otimes I_{D}$, where $M_{i,j}=\min(i,j)$, the spectrum of $K$ is equivalent to the spectrum of $M$ (just with all eigenvalues having multiplicity $D$). Therefore, we seek to find the spectrum of $M\in\mathbb{R}^{T\times T}$.
However, the spectrum of $M$ is known in the literature. For example, Theorem 2.1 of da Fonseca (2007) shows that if $T\geq 3$, then the eigenvalues $\{\lambda_{k}\}_{k=0}^{T-1}$ of $M$ are given by
$\lambda_{k}$ $=\frac{1}{2}\left(1-\cos\left(\frac{2k+1}{2T+1}\pi\right)\right)^{-1}$
$=\frac{1}{4}\left(\sin\left(\frac{2k+1}{2(2T+1)}\pi\right)\right)^{-2}$
where the second equality comes from the half-angle formula. We observe that the largest eigenvalue is therefore $\lambda_{0}$, and so the result follows after we take a square root. ∎
Appendix E Experimental Details and Additional Experiments